OpenStudy (anonymous):

Solve for x by rewriting both sides as powers of the same base: 3^x= (9^(x-1)) 27^(1-3x)

OpenStudy (anonymous):

there are three different bases in this equation... 3, 9 and 27... all three have their own powers.. but that really doesnt affect the processin this case... you want to change 3 9 and 27 into the same base.. meaning the same number raised to whatever power will give you the original number.. ex. 9 becomes 3^2... 27 becomes 3^3..\[3^{x} = 3^{2(x-1)}\times3^{3(1-3x)}\]

OpenStudy (anonymous):

then you can simplify the powers 2(x-1) and 3(1-3x)

OpenStudy (anonymous):

I get that but don't know how to solve this particular type of problem

OpenStudy (anonymous):

so its 2x-2 and 3-9x

OpenStudy (anonymous):

then join like terms -7x -1?

OpenStudy (anonymous):

yeah, those are the powers of the final answer

OpenStudy (anonymous):

if the problem just asks to change the numbers to the same base, then that answer is enough

OpenStudy (anonymous):

it says solve for x though

OpenStudy (anonymous):

oh.. woops didn't see that part

OpenStudy (anonymous):

thats where Im having trouble

OpenStudy (anonymous):

its cool

OpenStudy (anonymous):

once you get the same base, take the \[\log_{3} \] of both sides, which will leave you with x = 2x-2 + 3 - 9x

OpenStudy (anonymous):

remember when you multiply bases, you add the powers.. so when you take the log of both sides the base is removed and the powers become added

OpenStudy (anonymous):

ok, so x=-7x + 1 8x=1 x=1/8?

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yeah, that's right

OpenStudy (anonymous):

great, thanks man

OpenStudy (anonymous):

no problem, glad i could help

OpenStudy (anonymous):

keep posting, i'll keep answering for the next hour or so

OpenStudy (anonymous):

okay will do if I get stuck with another problem